You have a 5 question test with 4 multiple choice answers for each. What is your
probability of getting 3 correct answers?

Answer by stanbon(75887)

You have a 5 question test with 4 multiple choice answers for each. What is your
probability of getting 3 correct answers?
-----
Binomial Problem with p = 1/4 and n = 5
---
P(x=3) = 5C3(1/4)^3(3/4)^2 = binompdf(5,1/4,3)= 0.0879
=========================
Cheers,
Stan H.